# Exoplanets synchronization: Learning from Venus’ retrograde rotation

### v1

@@ 1 Introduction
This paper aims to use our knowledge of Venus’ rotation as an insight into what to expect for the rotation of Earth-like exoplanets in the habitable zone of a solar-type star. It collects various published results and compares them to hypothetical situations to improve our understanding of their meanings. It includes a new discussion of the existence and stability of stationary solutions under the joint actions of tidal and atmospheric torques (Auclair-Desrotour et al. 2017). When the eccentricity is neglected, these solutions might be synchronous or asynchronous (a.k.a. parasynchronous), stable or unstable (Correia and Laskar, 2003; Leconte et al., 2015). Two parameters indicating the ratios of the tidal and atmospheric torques are introduced, as well as a toy model. They allow us to study several different evolutionary paths for the rotation of exoplanets located in the habitable zone of solar-type stars. In these cases, the HZ is not too close to the star, and planets may harbor dynamically important atmospheres. The exploration of the model shows that the evolution leading to a retrograde rotation can happen at a slow pace and is not a low-probability event. The retrograde rotation of Venus may be the result of one smooth process of this kind, associated with the formation of its atmosphere, rather than a collisional one that has transferred a huge amount of negative angular momentum to the planet. The collisional processes often used to explain the retrograde rotation of Venus (McCord, 1968; Makarov and Goldin, 2023) are not the only possibility.
In section 2, the new full 3D version of the creep tide theory (Folonier et al. 2025) is used to study the evolution of the rotation of a hypothetical naked Venus in both the prograde and retrograde rotation cases. In particular, it is used to investigate how long the retrograde rotation can resist the action of the tidal torque. The precession of the planet’s equinoxes and the variations in the obliquity are also studied. In section 3, the torque due to the phase lead of the atmospheric bulge is introduced, and the main effects due to the joint actions of the atmospheric and tidal torques are investigated. In section 4, it is shown how the evolution of the atmosphere can cause the reversal of the planet’s rotation from prograde to retrograde. A toy model for the planet rotation provides evolutionary examples of this process. Section 5 collects the main conclusions of the report.
@@ 2 The naked Venus
Venus currently has a retrograde rotation with a period $243.0226\pm 0.0013$ days (Margot et al. 2021). We know that this rotation is strongly affected by the thermal bulges in the planet’s atmosphere. How would a retrograde rotation of this kind evolve should the planet be devoid of an atmosphere? We have investigated this topic using the new version of the creep tide theory valid for systems with a nonzero obliquity (Folonier et al., 2025).
@@ 2.1 Back to prograde
The main result is that the retrograde rotation of a naked Venus would be changed to prograde in a relatively short time (see Fig. 1). Using typical values for Venus’ parameters involved in this calculation (see Table 1), the result is an inversion in $\sim 700,000$ years, if the relaxation is $\gamma\sim 2\times 10^{-7}{\rm s}^{-1}$ (as the solid Earth cf. Ferraz-Mello, 2013). The tidal theory also indicates that at the inversion, we do not have a reversal of the rotation but a continuous migration of the rotation poles, which are displaced until the completion of the inversion. During the inversion interval, the angular momentum of the planet is very small (the rotation period grows to several thousand days). We do not know if the planet itself would turn upside down because the equations of the polhode are not included in the model. After the reversal of the rotation back to a prograde state, the rotation accelerates and is driven to a stable synchronization with the orbital period (see Fig. 1 Bottom).
Figure 1: Variation of the rotational elements of the planet before and after the rotation reversal: Top and Middle: Obliquity and Equinox before and after the rotation reversal. Bottom: Rotation period and Orbital period (blue line). Longitude of the First Equinoctial Point (intersection of the planet’s equator and orbit)
@@ 2.2 Precession
In the calculations leading to the results shown in Fig. 1, we have fixed the oblateness of the planet at $6.64\times 10^{-6}$. It is known that the oblatenesses of slow-rotating bodies (as Mercury and Venus) are much larger than predicted by the usual Maclaurin or Clairaut formulas. The model adopted here opts for the introduction of the oblateness by hand when the rotation velocity falls below a certain lead-off limit. This is done following the recipe: If $T_{\rm rot}\geq 33.528\$days, then $\epsilon_{\rm M}=6.42\times 10^{-6}$. The chosen lead-off period is the rotation period corresponding to the oblateness calculated from the $J_{2}$ value determined in recent space missions. The difference $0.22\times 10^{-6}$ between the two values mentioned above comes from the fact that the actual oblateness of the body is the sum of the rotational component and one small component due to the tidal deformation: $\epsilon_{\rm Z}=\epsilon_{\rm M}+\frac{1}{2}\epsilon_{\rm Jeans}$ (see Ferraz-Mello, 2015).
@@
Figure 2: Motion of the pole in the interval $660<t<720$ (kyr). Polar coordinates: Obliquity and Longitude of the First Equinoctial Point ($\Omega_{\rm Ipl}$). Dashed line: $\beta=90^{\circ}$.
@@ 2.3 Prograde rotation dynamics
Let us assume that the primordial rotation of Venus was prograde. How would this rotation be modified by tides in the absence of an atmosphere? The simulations show very different behaviors for fast and slow rotating planets. Fig. 3 shows (magenta lines) the results obtained assuming a solid Earth-like tidal relaxation ($\gamma=2\times 10^{-7}{\rm s}^{-1}$), and an initial rotation period of 10 days (terrestrial days). The almost synchronous stationary solution is reached in about 107 Myr  [^1] The stationary solution is slightly supersynchronous by an amount of the order of $e^{2}$. In the case of terrestrial planets, $T_{\rm pl}\sim T_{\rm orb}(1-6\gamma^{2}e^{2}/n^{2})$ (see Ferraz-Mello, 2013; Ferraz-Mello et al., 2025) which gives for a prograde Venus, a difference of the two periods $\Delta T/T\sim 10^{-4}$. With a nonzero eccentricity, synchronization can only be reached if the almost synchronous rotation entails the outcrop of an asymmetry in the mass distribution of the planet the torque of which counterbalances the small torque surplus due to eccentricity. The obliquity of the planet grows monotonically, but before reaching the stationary solution, it has a maximum and is then driven to zero in a few million years. Note that no obvious symmetry appears in the simulation as a cause of this behavior. This behavior seems very robust and does not depend critically on the initial conditions. Simulations with small variations in the initial conditions show only small, regular quantitative differences (see fig. 4). The chaotic behavior predicted by Laskar and Robutel (1993) does not affect the solutions because the secular perturbations due to the other planets of the system are not being considered. The equinoctial nodes precess initially with a period of $\sim 130$ kyr and grow following a quadratic time law to reach about 320 kyr in 100 Myr. When obliquity is driven to zero, the equinoctial nodes are no longer defined, and the circulation is converted to libration after a period of somewhat irregular behavior.
@@ 2.4 The rotation phase diagram
In the zero-obliquity version (a.k.a. coplanar version) of the tidal theory, the variation of the rotation of the planet is given by a first-order differential equation (Ferraz-Mello, 2013; Correia et al., 2014) of the form
@@ 3 The atmospheric torque
We have seen that the retrograde rotation of Venus could not survive for a long time, and the observed retrograde rotation is a clear clue to the existence of a torque acting in that direction. It has been attributed to a phase lead in the thermal deformation of the atmosphere (Gold and Soter, 1971; Ingersoll and Dobrovolskis, 1978)  [^2] For detailed discussions of this model, see Correia and Laskar(2003), Leconte et al.(2015), Auclair-Desrotour et al. (2017, 2024), and references therein.
Solar radiation raises the air temperatures in the afternoon and causes the air mass to flow to colder regions. In a first approximation, the atmosphere may be represented by an ellipsoid whose bulge is displaced in the direction contrary to the rotation of the body (see fig. 6). In this figure, the direction of the orbital motion does not matter. The rotation direction indicated is not absolute, but relative to the Sun’s direction. At one point on the planet’s surface, the tidal bulge is felt after noon. The tide lags the forcing tidal forces. In contrast, the atmospheric bulge leads the rotation. The gravitational effects of the Sun on the atmosphere introduce a torque in a direction opposite to that of the gravitational tidal torque.
@@
where $A^{\prime}$ is a positive coefficient  [^3] The coefficient $A^{\prime}$ depends on the mean motion ($n$), the radius ($R$), the moment of inertia ($C$), and also on the temperature at the surface, ($T_{0}$), on its specific heat at constant pressure, ($c_{p}$), and on the heat flux at the sub-solar point, ($\Phi_{0}$). It is such that $A^{\prime}\gamma^{\prime}=\frac{3\pi n^{2}R^{4}\Phi_{0}}{8c_{p}T_{0}C}.$ (5) This coefficient formally differs from the coefficient appearing in Ingersoll and Dobrovolskis(1978) by a factor 2. The division by 2 is here necessary to compensate the factofor 2 present in the expression of $\sin 2\sigma^{\prime}$. $A^{\prime}$ has the same dimension as $A$. and $\sigma^{\prime}$ is the lead of the atmospheric bulge defined by
$\tan\sigma^{\prime}=\frac{\nu}{\gamma^{\prime}}\equiv\nu\tau^{\prime}$ (6)
(Ingersoll and Dobrovolskis, 1978). Note that when $e=0$ equations (1) and (4) are formally equal. Since the given definitions of $\sigma$ and $\sigma^{\prime}$ are not univocal, we stress that in both cases they are defined by the principal value of the function, in the interval $(-\pi/2,+\pi/2)$. In the case of Venus, the maxima of the atmospheric pressure are recorded at 09:45 and 22:00 local solar time (see Auclair-Desrotour et al. 2024). This means $\sigma^{\prime}\sim 32^{\circ}$ .
@@ 3.1 Composite phase diagram
The parameter $b$ is akin to the ratio of the two torques. $b=0$ in the absence of atmospheric torque and it increases with the influence of the atmospheric torque. It plays a critical role on the existence and stability of stationary solutions.
@@ 3.2 Existence and stability of the asynchronous solutions
Equation (8) can be expressed as the quotient of two polynomials in $\nu$, and equation $\dot{\Omega}=0$ has the same roots as the symmetric third-degree polynomial in the numerator. One of these roots is $\nu_{0}=0$, and the other two are given by
@@
The condition for the existence of real roots is that the numerator and denominator are not simultaneously positive or negative (in which case we get $\nu_{0}^{2}<0$ and the two roots are imaginary). The conditions for the existence of the real roots $\pm\nu_{0}$ are:
@@
The analysis of the behavior of $\dot{\Omega}(\Omega)$ when $|\Omega|\rightarrow\infty$ allows us to conclude that, in case I, the synchronous root is unstable and the two asynchronous roots are stable (as shown in Fig. 7). In case II, the behavior is inverted, the synchronous root becomes stable and the two asynchronous roots become unstable.
Figure 8: Loci of the asynchronous stationary solutions for given values of $\Gamma$. These solutions are stable if they fall in the zone $b<1$ and unstable otherwise. If $|\nu_{0}/n|<1$ they are both prograde. If $|\nu_{0}/n|>1$, one is prograde and the other is retrograde.
Figure 9: Domains of existence of the asynchronous stationary solutions and vice versa, i stable, II unstable. Outside these domains, they are not real. In domains 0 and II, the synchronous stationary solution is stable. In other domains, it is unstable. The inserted sketches show how the roots are formed in each domain.
These stationary solutions are similar to those found by Auclair-Desrotour et al. (2017) using the classical Darwin tidal theory with Maxwell rheology.
@@ 3.3 Stability of the synchronous solutions
In domains I and II the synchronous solution is unstable when the asynchronous solutions are stable, and vice versa. In the other domains, the synchronous solution is the only real root. The stability is again deduced from the study of the leading term of the third-degree polynomial that defines the roots: $\dot{\Omega}\propto 2B(b-1)\nu_{0}^{3}+\&c.$ [^4] The function is of the kind y’=y.f(y). The stability of the root $y=0$ may be inferred from the study of a finite interval. The synchronous solution is stable when $b<1$ and unstable otherwise. This is consistent with the plots of the partial contributions of the tidal and atmospheric torques shown in Fig. 7.
@@
This result depends on the theory of tidal friction used. The Darwin theory with a Maxwell rheology (Auclair-Desrotour et al. 2017) gives results equal to those discussed above. However, if a constant lag or the lag given by Andrade’s rheology is plugged into Darwin’s theory, the asynchronous solutions come in pairs (Leconte et al. 2015), and the synchronous solution remains stable, even in the presence of stable asynchronous solutions.
@@ 4 The evolution towards a retrograde rotation
However, the impossibility mentioned above results from considering a frozen model. The inspection of Fig. 9 shows that, if we accept that the physical conditions are certainly not frozen, and the relative contribution of the tidal and atmospheric torques may have been different in the past, it is possible to have paths leading smoothly from a fast prograde rotation in the past to an asynchronous retrograde rotation in the present. We may, for example, consider an evolutive path starting in domain 0 (green area) of Fig. 9 and crossing the frontier between it and domain I. In domain 0, the rotation will evolve towards the synchronous solution and will reach it if enough time is allowed for the path segment inside domain O. When crossing the boundary towards domain I, the synchronous solution becomes unstable, and the rotation is forced to evolve towards one of the two new stable asynchronous solutions. If the rotation is synchronous before the crossing, there is no mathematical reason to privilege one of the two and we have equal probabilities of evolving towards $\mathsf{P}$ or $\mathsf{P}^{\prime}$. However, the increase in the moment of inertia due to the transfer of mass from the planet core to the atmosphere may create a bias in favor of retrograde rotations.
If the atmosphere of Venus is still changing, the current rotational state of Venus can change. For instance, if we assume that Venus’ rotation is currently trapped in a retrograde asynchronous stationary solution, we get $b=0.97489$, which is very close to the boundary $b=1$. A small change, say, in the surface temperature of Venus, may be such that $b$ increases to $b>1$. Hence, stable asynchronous stationary solutions cease to exist and the rotation of Venus will grow continuously.

### v2

@@ 1 Introduction
This paper aims to use our knowledge of Venus’ rotation as an insight into what to expect for the rotation of Earth-like exoplanets in the habitable zone of a solar-type star. It collects various published results and compares them to hypothetical situations to improve our understanding of their meanings. It includes a new discussion of the existence and stability of stationary solutions under the joint actions of tidal and atmospheric torques (Auclair-Desrotour et al. 2017). When the eccentricity is neglected, these solutions might be synchronous or asynchronous (a.k.a. parasynchronous), stable or unstable (Correia and Laskar, 2003; Leconte et al., 2015). Two parameters indicating the ratios of tidal and atmospheric torques are introduced, along with a toy model. They allow us to study several different evolutionary paths for the rotation of exoplanets located in the habitable zone of solar-type stars. In these cases, the HZ is not too close to the star, and planets may harbor dynamically important atmospheres. The exploration of the model shows that the evolution leading to a retrograde rotation can happen at a slow pace and is not a low-probability event. The retrograde rotation of Venus may be the result of one smooth process of this kind, associated with the formation of its atmosphere, rather than a collisional one that has transferred a huge amount of negative angular momentum to the planet. The collisional processes often used to explain the retrograde rotation of Venus (McCord, 1968; Makarov and Goldin, 2023) are not the only possibility. The smooth process proposed in this paper was first described by Correia et al. (2003a,b).
In section 2, the new full 3D version of the creep tide theory (Folonier et al. 2025) is used to study the evolution of the rotation of a hypothetical naked Venus in both the prograde and retrograde rotation cases. In particular, it is used to investigate how long the retrograde rotation can resist the action of the tidal torque. The precession of the planet’s equinoxes and the variations in the obliquity are also studied. In section 3, the torque due to the phase lead of the atmospheric bulge is introduced, and the main effects due to the joint actions of the atmospheric and tidal torques are investigated. In section 4, it is shown how the evolution of the atmosphere can cause the reversal of the planet’s rotation from direct to retrograde. A toy model for the planet rotation provides evolutionary examples of this process. Section 5 collects the main conclusions of the report.
@@ 2 The naked Venus
Venus currently has a retrograde rotation with a period of $243.0226\pm 0.0013$ days (Margot et al. 2021). We know that this rotation is strongly affected by the thermal bulges in the planet’s atmosphere. How would a retrograde rotation of this kind evolve should the planet be devoid of an atmosphere? We have investigated this topic using the new version of the creep tide theory valid for systems with a nonzero obliquity (Folonier et al., 2025).
@@ 2.1 Back to prograde
The main result is that the rotation of a naked Venus would be changed to prograde in a relatively short time (see Fig. 1). Using typical values for Venus’ parameters involved in this calculation (see Table 1), the result is an inversion in $\sim 700,000$ years, if the relaxation is $\gamma\sim 2\times 10^{-7}{\rm s}^{-1}$ (as the solid Earth cf. Ferraz-Mello, 2013). The tidal theory also indicates that at the inversion, we do not have a reversal of the rotation but a continuous migration of the rotation poles, which are displaced until the inversion is completed. During the inversion interval, the angular momentum of the planet is very small (the rotation period grows to several thousand days). We do not know if the planet itself would turn upside down because the equations of the polhode are not included in the model. After the reversal of the rotation back to a prograde state, the rotation accelerates and is driven to a stable synchronization with the orbital period (see Fig. 1 Bottom).
Figure 1: Variation of the rotational elements of the planet before and after the rotation reversal: Top and Middle: Obliquity and Equinox before and after the rotation reversal. Bottom: Rotation period (red) and Orbital period (blue). Longitude of the First Equinoctial Point (intersection of the planet’s equator and orbit).
@@ 2.2 Precession
In the calculations leading to the results shown in Fig. 1, we have fixed the oblateness of the planet at $6.64\times 10^{-6}$. We know that the oblatenesses of slow-rotating bodies (as Mercury and Venus) are much larger than predicted by the usual Maclaurin or Clairaut formulas. The model adopted here opts for the introduction of the oblateness by hand when the rotation velocity falls below a certain lead-off limit. This is done following the recipe: If $T_{\rm rot}\geq 33.528\$days, then $\epsilon_{\rm M}=6.42\times 10^{-6}$. The chosen lead-off period is the rotation period corresponding to the oblateness calculated from the $J_{2}$ value determined in recent space missions. The difference $0.22\times 10^{-6}$ between the two values mentioned above comes from the fact that the actual oblateness of the body is the sum of the rotational component and one small component due to the tidal deformation: $\epsilon_{\rm Z}=\epsilon_{\rm M}+\frac{1}{2}\epsilon_{\rm Jeans}$ (see Ferraz-Mello, 2015).
@@
Figure 2: Motion of the pole in the interval $660<t<720$ (kyr). Polar coordinates: Obliquity and Longitude of the First Equinoctial Point ($\Omega_{\rm Ipl}$). The line $\beta=90^{\circ}$ (dashed) is crossed from the outside in.
@@ 2.3 Prograde rotation dynamics
Let us assume that the primordial rotation of Venus was prograde. How would this rotation be modified by tides in the absence of an atmosphere? The simulations show very different behaviors for fast- and slow-rotating planets. Fig. 3 shows (magenta lines) the results obtained assuming a solid Earth-like tidal relaxation ($\gamma=2\times 10^{-7}{\rm s}^{-1}$), and an initial rotation period of 10 days (terrestrial days). The almost synchronous stationary solution is reached in about 107 Myr  [^1] The stationary solution is slightly supersynchronous by an amount of the order of $e^{2}$. In the case of terrestrial planets, $T_{\rm rot}\sim T_{\rm orb}(1-6\gamma^{2}e^{2}/n^{2})$ (see Ferraz-Mello, 2013; Ferraz-Mello et al., 2025) which gives for a prograde Venus, a difference of the two periods $\Delta T/T\sim 10^{-4}$. With a nonzero eccentricity, synchronization is only possible if the almost synchronous rotation causes an asymmetry in the planet’s mass distribution. The torque from this asymmetry must offset the small extra torque caused by the eccentricity. The obliquity of the planet grows monotonically, but before reaching the stationary solution, it has a maximum and is then driven to zero in a few million years. Note that no obvious symmetry appears in the simulation as a cause of this behavior. This behavior seems to be very robust and not critically dependent on the initial conditions. Simulations with small variations in the initial conditions show only small, regular quantitative differences (see fig. 4). The chaotic behavior predicted by Laskar and Robutel (1993) does not affect the solutions because the secular perturbations due to the other planets of the system are not being considered. The equinoctial nodes precess initially with a period of $\sim 130$ kyr and grow following a quadratic time law to reach about 320 kyr in 100 Myr. When obliquity is close to zero, the equinoctial nodes are no longer defined, and the circulation is converted to libration after a period of somewhat irregular behavior.
@@ 2.4 The rotation phase diagram
In the zero-obliquity version (a.k.a. coplanar version) of the tidal theory, the average variation of the rotation of the planet is given by a first-order differential equation (Ferraz-Mello, 2013; Correia et al., 2014) of the form
@@ 3 The atmospheric torque
We have seen that the retrograde rotation of Venus could not persist for a long time, and the observed retrograde rotation is a clear indication to the existence of a torque acting in that direction. It has been attributed to a phase lead in the thermal deformation of the atmosphere (Gold and Soter, 1971; Ingersoll and Dobrovolskis, 1978) [^2] For detailed discussions of this model, see Correia and Laskar(2003), Leconte et al.(2015), Auclair-Desrotour et al. (2017, 2024), and references therein.
Solar radiation raises the air temperatures in the afternoon and causes the air mass to flow to colder regions. In a first approximation, the atmosphere may be represented by an ellipsoid whose bulge is displaced in the direction contrary to the rotation of the body (see fig. 6). In this figure, the direction of the orbital motion does not matter. The rotation direction indicated is not absolute, but relative to the Sun’s direction. At one given point on the planet’s surface, the tidal bulge is felt after noon. The gravitational tide is delayed. In contrast, the atmospheric bulge leads the rotation. The gravitational effects of the Sun on the atmosphere introduce a torque in a direction opposite to that of the gravitational tide.
@@
where $A^{\prime}$ is a positive coefficient  [^3] The coefficient $A^{\prime}$ depends on the mean motion ($n$), the radius ($R$), the moment of inertia ($C$), and also on the temperature at the surface, ($T_{0}$), on its specific heat at constant pressure, ($c_{p}$), and on the heat flux at the sub-solar point, ($\Phi_{0}$). It is such that $A^{\prime}\gamma^{\prime}=\frac{3\pi n^{2}R^{4}\Phi_{0}}{8c_{p}T_{0}C}.$ (5) This coefficient formally differs from the coefficient appearing in Ingersoll and Dobrovolskis(1978) by a factor 2. The division by 2 is here necessary to compensate the factor 2 present in the expression of $\sin 2\sigma^{\prime}$. $A^{\prime}$ has the same dimension as $A$. and $\sigma^{\prime}$ is the lead of the atmospheric bulge defined by
$\tan\sigma^{\prime}=\nu\tau^{\prime}\equiv\frac{\nu}{\gamma^{\prime}}$ (6)
(Ingersoll and Dobrovolskis, 1978) where $\tau^{\prime}$ is a proportionality constant and $\gamma^{\prime}=1/\tau^{\prime}$. Note that when $e=0$, equations (1) and (4) are formally equal. Since the given definitions of $\sigma$ and $\sigma^{\prime}$ are not univocal, we stress that in both cases they are defined by the principal value of the function, in the interval $(-\pi/2,+\pi/2)$. In the case of Venus, the maxima of the atmospheric pressure are recorded at 09:45 and 22:00 local solar time (see Auclair-Desrotour et al. 2024). This means $\sigma^{\prime}\sim 32^{\circ}$ .
@@ 3.1 Composite phase diagram
The parameter $b$ is akin to the ratio of the two torques. It is $b=0$ in the absence of atmospheric torque, and it increases with the influence of the atmospheric torque. It plays a critical role in the existence and stability of stationary solutions.
@@ 3.2 Existence and stability of the asynchronous solutions
The bracket in Eqn. (8) can be expressed as the quotient of two polynomials in $\nu$, and equation $\dot{\Omega}=0$ has the same roots as the symmetric third-degree polynomial in the numerator. One of these roots is $\nu_{0}=0$, and the other two are given by
@@
The condition for the existence of real roots $\pm\nu_{0}$ is that the numerator and denominator are not simultaneously positive or negative (in which case we get $\nu_{0}^{2}<0$ and the two roots are imaginary):
@@
The analysis of the behavior of $\dot{\Omega}(\Omega)$ when $|\Omega|\rightarrow\infty$ allows us to conclude that, in case I, the synchronous root is unstable and the two asynchronous roots are stable (as shown in Fig. 7). In case II, the behavior is inverted, the synchronous root is stable, and the two asynchronous roots are unstable.
Figure 8: Loci of the asynchronous stationary solutions for given values of $\Gamma$. These solutions are stable if they fall in the zone $b<1$ and unstable otherwise. If $|\nu_{0}/n|<1$, they are both prograde. If $|\nu_{0}/n|>1$, one is prograde and the other is retrograde.
Figure 9: Domains of existence of the asynchronous stationary solutions: I stable, II unstable. Outside these domains, they are not real. In domains 0 and II, the synchronous stationary solution is stable. In other domains, it is unstable. The inserted sketches show how the roots are formed in each domain.
These stationary solutions are similar to those found by Auclair-Desrotour et al. (2017) using a Maxwell rheology.
@@ 3.3 Stability of the synchronous solutions
In domains I and II the synchronous solution is unstable when the asynchronous solutions are stable, and vice versa. In the other domains, the synchronous solution is the only real root. The stability is again deduced from the study of the leading term of the third-degree polynomial that defines the roots: $\dot{\Omega}={\rm const}\cdot(b-1)\nu^{3}+\&c.$ [^4] The function is of the kind y’=y.f(y). The stability of the root $y=0$ may be inferred from the study of a finite interval. The synchronous solution is stable when $b<1$ and unstable otherwise. This is consistent with the plots of the partial contributions of the tidal and atmospheric torques shown in Fig. 7.
@@ 4 The evolution towards a retrograde rotation
However, the impossibility mentioned above results from considering a frozen model. The inspection of Fig. 9 shows that, if we accept that the physical conditions are certainly not frozen, and the relative contribution of the tidal and atmospheric torques may have been different in the past, it is possible to have paths leading smoothly from a fast prograde rotation in the past to an asynchronous retrograde rotation in the present. We may, for example, consider an evolutive path starting in domain 0 (green area) of Fig. 9 and crossing the frontier between it and domain I. In domain 0, the rotation will evolve towards the synchronous solution and will reach it if enough time is allowed for the path segment inside domain 0. When crossing the boundary towards domain I, the synchronous solution becomes unstable, and the rotation is forced to evolve towards one of the two new stable asynchronous solutions. If the rotation is synchronous before the crossing, there is no mathematical reason to privilege one of the two and we have equal probabilities of evolving towards $\mathsf{P}$ or $\mathsf{P}^{\prime}$. However, the increase in the moment of inertia due to the transfer of mass from the planet core to the atmosphere may create a bias in favor of retrograde rotations.
If the atmosphere of Venus is still changing, the current rotational state of Venus can change. For instance, if we assume that Venus’ rotation is currently trapped in a retrograde asynchronous stationary solution, we get $b=0.97489$, which is very close to the boundary $b=1$. A small change, say, in the surface temperature of Venus, may be such that $b$ increases to $b>1$. Hence, stable asynchronous stationary solutions cease to exist, and the rotation of Venus will grow continuously.